# Revisiting Privacy, Utility, and Efficiency Trade-offs when Fine-Tuning Large Language Models

### v1

@@ Abstract
We study the inherent trade-offs in minimizing privacy risks and maximizing utility, while maintaining high computational efficiency, when fine-tuning large language models (LLMs). A number of recent works in privacy research have attempted to mitigate privacy risks posed by memorizing fine-tuning data by using differentially private training methods (e.g., DP), albeit at a significantly higher computational cost (inefficiency). In parallel, several works in systems research have focussed on developing (parameter) efficient fine-tuning methods (e.g., LoRA), but few works, if any, investigated whether such efficient methods enhance or diminish privacy risks. In this paper, we investigate this gap and arrive at a surprising conclusion: efficient fine-tuning methods like LoRA mitigate privacy-risks similar to private fine-tuning methods like DP. Our empirical finding directly contradicts prevailing wisdom that privacy and efficiency objectives are at odds during fine-tuning. Our finding is established by (a) carefully defining measures of privacy and utility that distinguish between memorizing sensitive and non-sensitive tokens in training and test datasets used in fine-tuning and (b) extensive evaluations using multiple open-source language models from Pythia, Gemma, and Llama families and different domain-specific datasets.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable abilities in solving varied natural language tasks [42], with a host of applications in multiple domains such as education [48], medical chatbot [46], AI assistant [10], etc. Their ability is acquired by first pre-training models on a large Web corpus[23], which results in general language understanding [6] and then fine-tuning the pre-trained models on domain-specifc data, which customizes their performance to specific tasks [55]. Two foundational challenges with pre-training/fine-tuning these models at scale are centered around enhancing privacy and efficiency, while retaining the models’ utility. The former is related to reducing the risk of LLMs leaking sensitive user information contained in the training data, while the latter is related to reducing the computational cost of training.
@@
A long line of recent research in the privacy community has focussed on methods to mitigate privacy risks when training LLMs [44, 56, 1, 51]. A notable example of such methods is differentially privacy (DP)[1]. Similarly, a flurry of recent research in the systems community has explored parameter-efficient fine-tuning (PEFT) methods in LLMs [14]. A notable example of this class of methods is low-rank adaptation (LoRA) [20]. However, no prior works, to the best of our knowledge, have investigated the privacy risks associated with the efficient training methods. So the central question driving our research here is: do efficient fine-tuning methods enhance or mitigate privacy risks during training?
@@
Our question above is motivated by the following high-level observations about DP and LoRA: methodologically, both DP and LoRA restrict the impact training examples can have on model parameters – DP deliberately through its noisy gradient update, and LoRA through low-rank adaptation. However, DP incurs significant additional computational overhead [11] compared to traditional fine-tuning, while LoRA significantly reduces the computational costs. So the answer to the above question can have significant consequences for achieving good privacy-efficiency-utility tradeoffs when fine-tuning. For instance, if LoRA mitigates privacy risks of training, then it suggests that this method can simultaneously achieve both privacy and efficiency objectives, contradicting the conventional wisdom drawn from DP literature that privacy comes at a computational cost. One of the key (surprising) findings of our work below lies in establishing that LoRA does indeed mitigate privacy risks.
When attempting to answer the above question, we encountered a more foundational question that has been not been well-addressed by the existing literature namely, how should one quantify the privacy risks associated with a fine-tuning method, so that it allows for a performance comparison across different methods? For instance, the claimed privacy benefits of the DP method are based on theoretical guarantees and not empirical evaluation, making it hard to compare it against privacy benefits of other fine-tuning methods.
Numerous recent works explored privacy concerns with LLMs that arise from their ability to memorize and recollect training data sequences containing sensitive personally identifiable information (PII), such as names, email addresses, phone numbers, credentials, etc. This sensitive information can leak from an LLM, either by accidental generation in response to benign requests to the model, or by deliberate extraction using adversarially crafted requests to the model[40, 38, 12, 22, 9, 41, 43].
However, surprisingly, existing privacy as well as utility metrics for LLMs ignore the distinction between sensitive and non-sensitive counterparts (tokens) in training data sequences [7, 40, 3]. [^1] A few works used this distinction to train models [44, 56], but it has not been applied when quantifying privacy risks or utility metrics. Sensitive data is often inherently random and less predictable, such as social security number or phone number. Non-sensitive data, on the other hand, is more predictable, e.g. there is only a small set of valid completions for the sentence “The dog chases the _". Consequently, the ability of LLMs to recollect or predict sensitive vs. non-sensitive data is very different (see Figure 1 for details). As we show in Section 3.2, ignoring this distinction leads to inaccurate measures of privacy and utility, and in particular, may have resulted in some prior works unintentionally exaggerating the threat of LLM’s memorizing training data[3, 9, 7].
Here, we propose and justify new measures for privacy and utility explicitly accounting for the differences in model’s performance for sensitive and non-sensitive data. For sensitive data, we are primarily interested in privacy, i.e. we want the loss on sensitive tokens to be as high as possible. Conversely, for non-sensitive data, we are interested in utility, i.e. we want the loss on non-sensitive tokens to be as low as possible.
@@
Figure 1: Sensitive and non-sensitive tokens have different predictability, measured as the recollection loss by pre-trained models in Figure 1(a) and fine-tuned models (recorded at epoch $5$) in Figure 1(b). The distinction motivates us to quantify privacy using sensitive token loss on training data (higher is better) and utility as non-sensitive token loss on test data (lower is better).
Contributions. We summarize our main findings below:
@@
2. Exaggerated estimates of privacy threat from LLM memorisation: We show that compared to our measure, existing measures significantly over-estimate privacy risks from LLM training. LLMs recollect non-sensitive tokens at a significantly higher level than sensitive tokens. So by not distinguishing between them, existing studies are over-estimating privacy threats.
3. Comparing privacy-utility-efficiency tradeoffs for three different fine-tuning methods: Our measures allow us to conduct a systematic and extensive empirical study of three different fine-tuning methods: full fine-tuning, DP, and LoRA, using models from three different LLM families, Pythia [4], Gemma [45] and Llama [47] over both real-world and synthetically-generated datasets. For each method and model, we evaluate privacy, utility (in terms of train and test loss respectively) and its impact on benchmark performance, and efficiency. Our comparative study yields several interesting insights: we find that – full fine-tuning results in poor utility-privacy tradeoffs; DP offers the best utility-privacy tradeoffs, but is computationally very expensive; LoRA is almost on par with DP in terms of utility and privacy, but is much more computationally efficient.
4. Feasibility of achieving all the three privacy, utility and efficiency objectives simultaneously: Our findings about LoRA performance suggest that it may be possible for methods to achieve all three objectives simultaneously. Our results challenge prevailing wisdom that enhancing privacy during training is necessarily more computationally expensive and call for investigating privacy benefits of existing and new parameter efficient fine-tuning methods.
@@ 2 Related Work
Privacy Attacks and Quantification. Privacy concerns in LLMs have gained attention in recent years [24, 49], particularly the possibility of exposing data via membership inference attacks (MIA) [40, 38, 12, 22] and training data extraction attacks [9, 41, 43]. While membership inference attacks have the goal of identifying if a certain datapoint was in the training dataset using model confidence scores, training data extraction attacks try to extract specific parts of data from the training dataset using different prompting strategies. A recent study [54] underscores that membership inference attacks may not be reliable due to their practical limitations of requiring knowledge of the entire training data, thus proving data extraction attacks to be more viable.
@@
Existing works have quantified privacy attacks in large language models especially through the lens of memorization. For example, the authors in [7] assess memorization by quantifying how closely LLM generation matches the exact training data phrases when prompted with tailored prefixes. The metric of ‘exposure’ is introduced in [8] to evaluate a model’s vulnerability when exposed to data that was artificially introduced into the dataset (also known as "canaries") several times during training. This metric has been widely adopted in subsequent research [44, 56] as a measure of privacy. However, the metric’s formulation is limited by its reliance on assumptions about the surrounding knowledge of other canaries, which do not hold in practical scenarios.
@@
Intuitively, privacy evaluation should be more related to sensitive data seen by the model during its training. Some of the existing studies  [25, 36, 56, 44] focusing on privacy attacks in LLMs may consider sensitive and non-sensitive counterparts in the data explicitly during training, but they do not retain this distinction during privacy quantification. We believe (and later show) that this lack of distinction can lead to an inaccurate (sometimes, overestimated) assessment of privacy threats. In our paper, we carefully make the distinction between sensitive and non-sensitive data and propose a revised quantification.
@@
Privacy-Utility Tradeoffs. To mitigate privacy leakage, differential privacy (DP) measures have been proposed, which add theoretical privacy guarantees to the training process [1]. The key idea of DP is to clip the gradients of each datapoint and add noise to the clipped gradients in every iteration. Its primary goal is to reduce the influence of individual datapoints in the training procedure thus preventing its leakage during inference.
@@
Existing work utilising DP [51] has shown a clear trade-off between privacy and utility where privacy is measured in terms of theoretical guarantee and utility in terms of overall performance on the end task. The authors in [5] also question the notion of privacy considered in these privacy preserving techniques. Another line of work from [44] and [56] distinguishes between sensitive and non-sensitive data using techniques like regex and redaction and proposes a customised training strategy using DP. However, their evaluation measure also relies on theoretical guarantees for privacy and overall performance for utility. The question that arises here is whether this measure of quantification of privacy and utility is sufficient in the context of large language models, or whether it needs to be more nuanced.
@@
Additionally, to ensure privacy guarantees, differential privacy requires higher computational resources (time and memory) during training. For example, compared to vanilla SGD, DP-SGD may incur up to $20$x training time, which is often a bottleneck in resource-intensive tasks [11]. Therefore, the conventional wisdom is that privacy comes at the cost of efficiency. This leads us to probe ways in which the above limitation can be prevented.
@@
Utility-Efficiency Tradeoffs. Fine-tuning a pre-trained LLM on a task specific context has been used in several applications like medical chatbots, AI assistants, etc. However, full fine-tuning of large language models is expensive as it involves updating all the parameters. Recent work has focused on reducing the training cost while maintaining utility, with the introduction of parameter efficient fine-tuning (PEFT) techniques [14]. Well-known PEFT methods include adapter-based fine-tuning [19, 27, 57, 15], soft prompt-based fine-tuning [30, 29, 33, 34, 28], and parameterized fine-tuning [20, 32]. Among different PEFT methods, the Low-Rank Adaptation method, called LoRA [20] has emerged as one of the most widely used methods. LoRA updates fewer parameters in the model via low-rank approximation, providing computational efficiency with a relatively low cost to utility.
@@
In recent work on private fine-tuning [35, 50, 37], the authors combine LoRA with DP to reduce the additional computational overhead induced by differential privacy. In such a context, we ask whether DP is the only method towards ensuring privacy. LoRA’s training procedure of updating fewer parameters can be thought of as analogous to the noisy gradient update in DP. This leads us to the question of whether LoRA has any privacy benefits besides having control over utility-efficiency tradeoffs.
@@ 3.1 Rethinking Privacy and Utility
Existing studies at the intersection of privacy and natural language processing [56, 44, 31, 52] seek to enhance privacy while maintaining model utility. Utility is generally assessed based on model performance, such as loss, accuracy, or perplexity across the entire test dataset. Privacy is evaluated in terms of performance measures on the entire training dataset or theoretical guarantees in differential privacy (DP).
Natural language text may contain both sensitive and non-sensitive words, referred to as entities. For example, sensitive entities include names, addresses, phone numbers, order IDs, and other personally identifiable information. In contrast, non-sensitive entities generally involve semantic and/or syntactic completions following predictable patterns in language generation tasks. Informally, sensitive entities are drawn from a large search space (e.g., a random sequence of digits), resulting in high entropy and low predictability. In contrast, non-sensitive entities are more restricted in their occurrences (e.g., a subject is typically followed by a verb), leading to low entropy and high predictability. Several studies [3, 44, 56] distinguish between sensitive and non-sensitive entities in their proposed privacy leakage mitigation methods. However, the distinction is not leveraged in the quantification of privacy and utility, which is essential for a granular evaluation as discussed next.
@@ I. Analyzing privacy and utility while fine-tuning an LLM
Results. In Figure 2, we demonstrate measures of privacy (left column) and utility (right column) while fine-tuning three LLM models on Customersim dataset [44] (experimental details are provided at the end of this section). In particular, we show training loss on the left column and test loss on the right column. Importantly, we separately compute the loss for both sensitive and non-sensitive tokens in both training and test datasets. Intuitively, a higher loss denotes more privacy and less utility.
Privacy is overestimated in the existing measure. In Figure 2(a), we compute privacy using our measure, as well as the existing one. The existing measure of privacy considers all tokens in the training data, where low training loss denotes less privacy, while our measure considers only the sensitive tokens in the training data. Using our measure, a notable disparity emerges: sensitive tokens exhibit significantly higher loss than non-sensitive ones, particularly in the initial training epochs, as sensitive tokens are less predictable. This eventually indicates that the loss over all tokens (existing measure) would be much lower initially than the loss over only sensitive tokens (our measure), thus overestimating privacy threats much earlier. Similar trends are observed for other models in Figures 2(c) and 2(e).
Utility is underestimated in the existing measure. In Figure 2(b), we compute utility using our measure and the prevailing one. The existing utility measure is related to the test loss of all tokens, where lower test loss indicates better utility. We can observe that our measure that considers the test loss on only non-sensitive tokens provides better utility than the existing measure. Similar trends are observed for other models in Figures 2(d) and 2(f). Existing measure would lead to better utility at earlier epochs compared to our measure i.e. one would select checkpoints at epochs $4$, $2$ and $2$ using existing measures for Pythia, Gemma and Llama2 respectively and at epochs $5$, $6$, and $6$ using our measure. Thus, utility is underestimated in the existing measure. Figure 18 in Appendix 14 shows the change more distinctly in log-scale. For the same fine-tuning scenario, our measure certifies better utility than the existing measure, which unintentionally considers the loss of sensitive tokens that are rare. The key to our findings is how we annotate sensitive and non-sensitive tokens in real-world datasets. Next, we expand our experimental details leading to our results.
@@
Experimental setup and methodology. We perform our analysis on two datasets: CustomerSim [44], a simulated dialog dataset for conversation generation and SynBio (originally called PII) [18], an LLM generated dataset representing student biographies containing personal identifiable information. Table 1 shows some excerpts from the datasets. We use three open-source models during evaluation: Pythia 1B [4], Gemma 2B[45], and Llama2 7B [47].
We leverage two tools for annotating sensitive information in a given text: Presidio [39], which helps in identification of private entities in text, and GPT-4 [2], which is provided with a particular prompt for returning the annotated portions. An example of such a prompt for annotating samples is provided in Appendix 11.
@@
To summarize, the degree of difference in computing privacy and utility using our measure and the existing one depends on the ratio of sensitive to non-sensitive tokens. A higher ratio would result in a higher difference in the measure, and vice versa. Considering the distinction between sensitive and non-sensitive tokens, we show that the existing measure can both exaggerate privacy threats and underestimate the utility in LLMs. In this context, we re-examine a prior study [3] to better support our claim that reported privacy threats are exaggerated.
@@
Table 2: Examples of memorized sequences from [3], often containing predictable and non-sensitive patterns, like mathematical series and licensing text.
@@
II. Examining memorized sequences from [3]:
We consider a case study to analyze the reported memorized strings by [3], Our goal is to examine whether the memorized strings contain sensitive information or mere syntactic and semantic patterns.
Experimental setup. The authors in [3] considered the task of predicting whether a model memorizes specific training data points from the Pile dataset [13], which is used to train base LLM models. Among published memorized strings, we randomly choose $5,000$ strings from the pythia-1b-dup split [3]. A representative list of memorized strings is in Table 2, where the strings often follow syntactic and semantic patterns, such as completion of mathematical series, code snippets, licensing agreements, etc. Therefore, our hypothesis is that most of the memorized strings contain a great amount of non-sensitive and highly predictable tokens. To validate our hypothesis, we query for the source of memorized strings with respect to the training dataset, Pile, which aggregates data from multiple sources such as Pile-Cc, OpenWebText, ArXiv, etc. We leverage GPT-4 model to accomplish our task – given a memorized string, we ask for the source of the string from the list of Pile sections. The prompt template for GPT-4 is the following.
{mdframed}
[backgroundcolor=lavender, linewidth=0pt] You are provided with the following text: {memorized–sequence}.

### v2

@@ Abstract
We study the inherent trade-offs in minimizing privacy risks and maximizing utility, while maintaining high computational efficiency, when fine-tuning large language models (LLMs). A number of recent works in privacy research have attempted to mitigate privacy risks posed by memorizing fine-tuning data by using differentially private training methods (e.g., DP-SGD), albeit at a significantly higher computational cost (inefficiency). In parallel, several works in systems research have focussed on developing (parameter) efficient fine-tuning methods (e.g., LoRA), but few works, if any, investigated whether such efficient methods enhance or diminish privacy risks. In this paper, we investigate this gap and arrive at a surprising conclusion: efficient fine-tuning methods like LoRA mitigate privacy-risks similar to private fine-tuning methods like DP-SGD. Our empirical finding directly contradicts prevailing wisdom that privacy and efficiency objectives are at odds during fine-tuning. Our finding is established by (a) carefully defining measures of privacy and utility that distinguish between memorizing sensitive and non-sensitive tokens in training and test datasets used in fine-tuning and (b) extensive evaluations using multiple open-source language models from Pythia, Gemma, Llama, and Qwen families and different domain-specific datasets.
@@ 1 Introduction
Large language models (LLMs) have shown proficiency across diverse natural language tasks [naveed2023comprehensive], finding applications in education [wang2024large], medical chatbots [thirunavukarasu2023large], and AI assistants [dong2023towards]. Their capabilities stem from a two-phase process: (a) pre-training on extensive web data [kaplan2020scaling] to develop general language understanding [brown2020language], (b) fine-tuning on domain-specific data for specific tasks [zhang2023instruction]. In both phases, the key challenges involve enhancing privacy and efficiency while maintaining the models’ utility. Privacy is related to reducing the risk of LLMs leaking sensitive user information contained in the training data, efficiency is related to reducing the computational cost of training, while utility is related to generative performance over test data.
@@
A long line of recent research in the privacy community has focussed on methods to mitigate privacy risks when training LLMs [shi-etal-2022-selective, zhao-etal-2022-provably, 10.1145/2976749.2978318, yu2022differentiallyprivatefinetuninglanguage]. A notable example of such methods is differential privacy based stochastic gradient descent (DP-SGD)[10.1145/2976749.2978318]. Simultaneously, a flurry of recent research in the systems community has explored parameter-efficient fine-tuning in LLMs [han2024parameterefficientfinetuninglargemodels]. A notable example of this class of methods is low-rank adaptation (LoRA) [hu2022lora]. However, no prior works, to the best of our knowledge, have investigated the privacy risks associated with efficient training methods.
@@
The central question driving our research here is: do efficient fine-tuning methods enhance or mitigate privacy risks during training? As DP-SGD incurs significant additional computational overhead [dupuy2022efficient], while LoRA significantly reduces the computational costs, the answer to the above question can have significant consequences for achieving good privacy-efficiency-utility tradeoffs when fine-tuning. For instance, if LoRA mitigates privacy risks of training that would suggest that it can simultaneously achieve both privacy and efficiency objectives, contradicting the conventional wisdom drawn from DP literature that privacy comes at a computational cost. A key (surprising) finding of our work lies in establishing that LoRA does indeed mitigate privacy risks.
A conjecture that might explain our finding is rooted in the following high-level observations about DP-SGD and LoRA: methodologically, both DP-SGD and LoRA restrict the impact that training examples can have on model parameters – DP-SGD deliberately through its noisy gradient update, and LoRA through low-rank adaptation. We formalize this intuition in this work.
When attempting to answer the above question, we encountered a more foundational question: how should one quantify such privacy risks associated with a fine-tuning method, so that it allows for a performance comparison across different methods? Numerous studies have highlighted privacy risks in LLMs due to their tendency to memorize and regurgitate training data containing sensitive personally identifiable information (PII) such as names, emails, and credentials  [mattern2023membership, fu2023practical, kaneko2024sampling, carlini2021extracting, mireshghallah-etal-2022-empirical, panda2024teach]. A natural way to account for privacy risks from memorization might be to measure loss on recollecting tokens in training data sequences.
However, we find that LLMs exhibit very different losses in recollecting sensitive vs. non-sensitive tokens in training data (see Figure 1). This difference is due to inherent randomness and unpredictability of sensitive data (e.g., phone numbers, SSNs) compared to non-sensitive data (e.g., “The dog chases the _”), which is often more structured and predictable. Consequently, we propose a new privacy measure that explicitly account for this difference: aiming for high loss on sensitive tokens from training data.
@@
Figure 1: Sensitive and non-sensitive tokens have different predictability, measured as the recollection loss by pre-trained models (Figure 1(a)) and fine-tuned models (Figure 1(b)). This distinction motivates to quantify privacy using sensitive tokens from the training data.
Our contributions and findings. The primary contribution of this work is to explore if efficient fine-tuning methods inherently mitigate privacy risks during training. To the best of our knowledge, no prior works have investigated these privacy risks associated with efficient training methods. We summarize our main contributions and findings as follows:
@@
2. Comparing privacy-utility-efficiency tradeoffs for three different fine-tuning methods: Our measures allow us to conduct a systematic and extensive empirical study of three different fine-tuning methods: full fine-tuning (FFT), DP-SGD, and LoRA, using models from four different LLM families, Pythia [biderman2023pythia], Gemma [team2024gemma], Llama [touvron2023llama] and Qwen [yang2024qwen2] over two datasets. Our comparative study yields several interesting insights: we find that – FFT results in poor utility-privacy tradeoffs; DP offers reasonable utility-privacy tradeoffs, but is computationally very expensive; LoRA is almost on par with DP in terms of privacy with good privacy-utility tradeoffs, but is more computationally efficient. To re-confirm our results, we evaluate the three methods using existing privacy measures—privacy loss [10.1145/2976749.2978318] and canary exposure [carlini2019secret]. We also show formally that the effect of fine-tuning a model using LoRA and DP on a datapoint is analogous, leading to privacy benefits.
3. Feasibility of achieving all the three privacy, utility and efficiency objectives simultaneously: Our findings about LoRA performance challenge prevailing wisdom that enhancing privacy during training is more computationally expensive. This calls for investigating privacy benefits of existing and new parameter efficient fine-tuning methods.
@@ 2 Related Work
Privacy Attacks and Quantification. Privacy concerns in LLMs have gained attention in recent years [KIBRIYA2024109698, YAO2024100211], particularly the possibility of exposing data via membership inference attacks (MIA) [mireshghallah-etal-2022-quantifying, mattern2023membership, fu2023practical, kaneko2024sampling] and training data extraction attacks [carlini2021extracting, mireshghallah-etal-2022-empirical, panda2024teach]. While membership inference attacks have the goal of identifying if a certain datapoint was in the training dataset using model confidence scores, training data extraction attacks try to extract specific parts of data from the training dataset using different prompting strategies. A recent study [zhang2024membershipinferenceattacksprove] underscores that membership inference attacks may not be reliable due to their practical limitations of requiring knowledge of the entire training data, thus proving data extraction attacks to be more viable.
@@
Existing works have quantified privacy attacks in large language models especially through the lens of memorization. For example, the authors in [carlini2023quantifying] assess memorization by quantifying how closely LLM generation matches the exact training data phrases when prompted with tailored prefixes. The metric of ‘exposure’ is introduced in [10.5555/3361338.3361358] to evaluate a model’s vulnerability when exposed to data that was artificially introduced into the dataset (also known as "canaries") several times during training. This metric has been widely adopted in subsequent research [shi-etal-2022-selective, zhao-etal-2022-provably] as a measure of privacy. However, the metric’s formulation is limited by its reliance on assumptions about the surrounding knowledge of other canaries, which do not hold in practical scenarios.
@@
Intuitively, privacy evaluation should be more related to sensitive data seen by the model during its training. Some of the existing studies  [10.5555/3666122.3667033, 10179300, zhao-etal-2022-provably, shi-etal-2022-selective] focusing on privacy attacks in LLMs may consider sensitive and non-sensitive counterparts in the data explicitly during training, but they do not retain this distinction during privacy quantification. We believe (and later show) that this lack of distinction can lead to an inaccurate (sometimes, overestimated) assessment of privacy threats. In our paper, we carefully make the distinction between sensitive and non-sensitive data and propose a revised quantification.
@@
Privacy-Utility Tradeoffs. To mitigate privacy leakage, differential privacy (DP-SGD-SGD) measures have been proposed, which add theoretical privacy guarantees to the training process [10.1145/2976749.2978318]. The key idea of DP-SGD-SGD is to clip the gradients of each datapoint and add noise to the clipped gradients in every iteration. Its primary goal is to reduce the influence of individual datapoints in the training procedure thus preventing its leakage during inference.
@@
Existing work utilising DP-SGD [yu2022differentiallyprivatefinetuninglanguage] has shown a clear trade-off between privacy and utility where privacy is measured in terms of theoretical guarantee and utility in terms of overall performance on the end task. The authors in [10.1145/3531146.3534642] also question the notion of privacy considered in these privacy preserving techniques. Another line of work from [shi-etal-2022-selective] and [zhao-etal-2022-provably] distinguishes between sensitive and non-sensitive data using techniques like regex and redaction and proposes a customised training strategy using DP-SGD. However, their evaluation measure also relies on theoretical guarantees for privacy and overall performance for utility. The question that arises here is whether this measure of quantification of privacy and utility is sufficient in the context of large language models, or whether it needs to be more nuanced.
@@
Additionally, to ensure privacy guarantees, differential privacy requires higher computational resources (time and memory) during training. For example, compared to vanilla SGD, DP-SGD-SGD may incur up to 2020x training time, which is often a bottleneck in resource-intensive tasks [dupuy2022efficient]. Therefore, the conventional wisdom is that privacy comes at the cost of efficiency. This leads us to probe ways in which the above limitation can be prevented.
@@
Utility-Efficiency Tradeoffs. Fine-tuning a pre-trained LLM on a task specific context has been used in several applications like medical chatbots, AI assistants, etc. However, full fine-tuning of large language models is expensive as it involves updating all the parameters. Recent work has focused on reducing the training cost while maintaining utility, with the introduction of parameter efficient fine-tuning (PEFT) techniques [han2024parameterefficientfinetuninglargemodels]. Well-known PEFT methods include adapter-based fine-tuning [houlsby2019parameter, lei2023conditional, zhu2021counter, he2021towards], soft prompt-based fine-tuning [li2021prefix, li2023prefix, liu2021p, liu2024gpt, lester2021power], and parameterized fine-tuning [hu2022lora, liu2024dora]. Among different PEFT methods, the Low-Rank Adaptation method, called LoRA [hu2022lora] has emerged as one of the most widely used methods. LoRA updates fewer parameters in the model via low-rank approximation, providing computational efficiency with a relatively low cost to utility.
@@
In recent work on private fine-tuning [liu2023differentially, yu2021differentially, ma2024efficient], the authors combine LoRA with DP-SGD to reduce the additional computational overhead induced by differential privacy. In such a context, we ask whether DP-SGD is the only method towards ensuring privacy. LoRA’s training procedure of updating fewer parameters can be thought of as analogous to the noisy gradient update in DP-SGD. This leads us to the question of whether LoRA has any privacy benefits besides having control over utility-efficiency tradeoffs.
@@ 3.1 Rethinking Privacy and Utility
Existing studies at the intersection of privacy and natural language processing [zhao-etal-2022-provably, shi-etal-2022-selective, li2022large, yu2022differentially] seek to enhance privacy while maintaining model utility. Utility is generally assessed based on model performance, such as loss, accuracy, or perplexity across the entire test dataset. Privacy is evaluated in terms of performance measures on the entire training dataset or theoretical guarantees in differential privacy (DP-SGD).
Natural language text may contain both sensitive and non-sensitive words, referred to as entities. For example, sensitive entities include names, addresses, phone numbers, order IDs, and other personally identifiable information. In contrast, non-sensitive entities generally involve semantic and/or syntactic completions following predictable patterns in language generation tasks. Informally, sensitive entities are drawn from a large search space (e.g., a random sequence of digits), resulting in high entropy and low predictability. In contrast, non-sensitive entities are more restricted in their occurrences (e.g., a subject is typically followed by a verb), leading to low entropy and high predictability. Several studies [biderman2024emergent, shi-etal-2022-selective, zhao-etal-2022-provably] distinguish between sensitive and non-sensitive entities in their proposed privacy leakage mitigation methods. However, the distinction is not leveraged in the quantification of privacy and utility, which is essential for a granular evaluation as discussed next.
@@ 3.2 Why do we distinguish between sensitive and non-sensitive entities?
(g) Privacy measure in Qwen2.5
(h) Utility measure in Qwen2.5
@@ I. Analyzing privacy and utility while fine-tuning an LLM
Results. In Figure 2, we demonstrate measures of privacy (left column) and utility (right column) while fine-tuning three LLM models on Customersim dataset [shi-etal-2022-selective] (experimental details are provided at the end of this section). In particular, we show training loss on the left column and test loss on the right column. Importantly, we separately compute the loss for both sensitive and non-sensitive tokens in both training and test datasets. Intuitively, a higher loss denotes more privacy and less utility.
Privacy is overestimated in the existing measure. In Figure 2(a), we compute privacy using our measure, as well as the existing one. The existing measure of privacy considers all tokens in the training data, where low training loss denotes less privacy, while our measure considers only the sensitive tokens in the training data. Using our measure, a notable disparity emerges: sensitive tokens exhibit significantly higher loss than non-sensitive ones, particularly in the initial training epochs, as sensitive tokens are less predictable. This eventually indicates that the loss over all tokens (existing measure) would be much lower initially than the loss over only sensitive tokens (our measure), thus overestimating privacy threats much earlier. Similar trends are observed for other models in Figures 2(c), 2(e), and 2(g).
Utility is underestimated in the existing measure. In Figure 2(b), we compute utility using our measure and the prevailing one. The existing utility measure is related to the test loss of all tokens, where lower test loss indicates better utility. We can observe that our measure that considers the test loss on only non-sensitive tokens provides better utility than the existing measure. Similar trends are observed for other models in Figures 2(d), 2(f), and 2(h). Existing measure would lead to better utility at earlier epochs compared to our measure i.e. one would select checkpoints at epochs $4$, 22, 22, and $2$ using existing measures for Pythia, Gemma, Llama2, and Qwen2.5 respectively and at epochs $5$, 66, 55, and $6$ using our measure. Thus, utility is underestimated in the existing measure. For the same fine-tuning scenario, our measure certifies better utility than the existing measure, which unintentionally considers the loss of sensitive tokens that are rare. The key to our findings is how we annotate sensitive and non-sensitive tokens in real-world datasets. Next, we expand our experimental details leading to our results.
@@
Experimental setup and methodology. We perform our analysis on two datasets: CustomerSim [shi-etal-2022-selective], a simulated dialog dataset for conversation generation and SynBio (originally called PII) [pii], an LLM generated dataset representing student biographies containing personal identifiable information. Table 1 shows some excerpts from the datasets. We use four open-source models during evaluation: Pythia-1B [biderman2023pythia], Gemma-2B [team2024gemma], Llama2-7B [touvron2023llama], and Qwen2.5-7B [yang2024qwen2]
We leverage two tools for annotating sensitive information in a given text: Presidio [MsPresidio], which helps in identification of private entities in text, and GPT-4 [achiam2023gpt], which is provided with a particular prompt for returning the annotated portions. An example of such a prompt for annotating samples is provided in Appendix 14.
@@
To summarize, the degree of difference in computing privacy and utility using our measure and the existing one depends on the ratio of sensitive to non-sensitive tokens. A higher ratio would result in a higher difference in the measure, and vice versa. Considering the distinction between sensitive and non-sensitive tokens, we show that the existing measure can both exaggerate privacy threats and underestimate the utility in LLMs. In this context, we re-examine a prior study [biderman2024emergent] to better support our claim that reported privacy threats are exaggerated.
@@
Table 2: Examples of memorized sequences from [biderman2024emergent], often containing predictable and non-sensitive patterns, like mathematical series and licensing text.
@@
II. Examining memorized sequences from [biderman2024emergent]:
We consider a case study to analyze the reported memorized strings by [biderman2024emergent], Our goal is to examine whether the memorized strings contain sensitive information or mere syntactic and semantic patterns.
Experimental setup. The authors in [biderman2024emergent] considered the task of predicting whether a model memorizes specific training data points from the Pile dataset [gao2020pile], which is used to train base LLM models. Among published memorized strings, we randomly choose $5,000$ strings from the pythia-1b-dup split [biderman2024emergent]. A representative list of memorized strings is in Table 2, where the strings often follow syntactic and semantic patterns, such as completion of mathematical series, code snippets, licensing agreements, etc. Therefore, our hypothesis is that most of the memorized strings contain a great amount of non-sensitive and highly predictable tokens. To validate our hypothesis, we query for the source of memorized strings with respect to the training dataset, Pile, which aggregates data from multiple sources such as Pile-Cc, OpenWebText, ArXiv, etc. We leverage GPT-4 model to accomplish our task – given a memorized string, we ask for the source of the string from the list of Pile sections. The prompt template for GPT-4 is the following.
You are provided with the following text: {memorized–sequence}.